OpenStudy (anonymous):

“The sum of 4 and 6 is the same as 2 times the sum of 2 and 3.” Which property of numbers did Frank apply to write the statement? Answer addition property of equality associative property of addition commutative property of addition distributive property of addition

OpenStudy (cwrw238):

4 + 6 = 2(2 + 3) distributive property of addition

OpenStudy (anonymous):

4+6=2(2+3)

OpenStudy (anonymous):

ight thanks!

OpenStudy (cwrw238):

yw